As a block of ice is melted (to 4 °C). Its volume.
Option A. remains the same.
Option B. decreases.
Option C. increases.


Which one is true
1
Expert's answer
2019-09-04T09:14:51-0400

It's volume decreases.

Density of is is less than it is of water:

"d(ice)≈0.916\frac{g}{cm^3}"

Water has maximum density at "4\u00b0C"

"d(water)= 1\frac{g}{cm^3}"


If we have 1 g of ice its volume is

"V(ice)=\frac{m}{d}= \frac{1}{0.916}=1.092 cm^3"

Volume of 1 g of melted ice is volume of water

"V(water)= \frac{1}{1}=1 cm^3"


So, we see that when ice is melted to water at 4oC density decreases.